# Navigating the Metrics Maze: Reconciling Score Magnitudes and Accuracies

### v1

@@ Abstract
Ten years ago a single metric, BLEU, governed progress in machine translation research. For better or worse, there is no such consensus today, and consequently it is difficult for researchers to develop and retain the kinds of heuristic intuitions about metric deltas that drove earlier research and deployment decisions. This paper investigates the “dynamic range” of a number of modern metrics in an effort to provide a collective understanding of the meaning of differences in scores both within and among metrics; in other words, we ask what point difference in metric is required between two systems for humans to notice? We conduct our evaluation on a new large dataset, ToShip23, using it to discover deltas at which metrics achieve system-level differences that are meaningful to humans, which we measure by pairwise system accuracy. We additionally show that this method of establishing delta-accuracy is more stable than the standard use of statistical p-values in regards to testset size. Where data size permits, we also explore the effect of metric deltas and accuracy across finer-grained features such as translation direction, domain, and system closeness.
@@ 1 Introduction
A decade ago, the BLEU metric served as the default metric for machine translation evaluation. It was not without its criticisms (Hovy and Ravichandran, 2003; Callison-Burch et al., 2006; Belz and Reiter, 2006) or compelling alternatives (Banerjee and Lavie, 2005; Popović, 2015), but a combination of adequate performance, robustness to new languages, simplicity and understandability, and inertia helped it retain this position. This is no longer the case. BLEU’s deficiencies quickly became apparent as deep learning approaches to machine translation replaced the earlier symbolic paradigms (Mathur et al., 2020a). Today, a number of metrics—themselves deep-learning based—compete in an ecosystem where there is no longer any dominant, default metric.
This situation creates a problem for researchers working to keep abreast of developments in the field. Different metrics, including different models within the same metric family, have different “dynamic ranges”, i.e., the range of scores one can expect to see. Furthermore, the “metric delta”, i.e., the score difference signifying a meaningful change in performance between two systems, also varies across metrics. It is perhaps understandable that some practitioners therefore continue to use BLEU, as well, if only to ground their understanding.
@@ 2 Experimental Setup
We perform experiments related to evaluation of MT outputs based on a proprietary dataset ToShip23 which is of a magnitude larger than any publicly available data and enables more fine grained glimpse into the metrics behaviour. The dataset is an extended version of ToShip21 dataset (Kocmi et al., 2021) with details described in Appendix B. We also use data from the annual WMT evaluation campaigns to validate our results, specifically the metrics shared task (Freitag et al., 2022, 2023), to make results replicable. We only use MQM (Freitag et al., 2021a) and DA+SQM (Kocmi et al., 2022) subset of human evaluated systems because reference-based DA (Bojar et al., 2016) is suboptimal for the evaluation of modern MT systems (Freitag et al., 2022). See Table 1 for an overview of dataset sizes.
@@
We evaluate the most frequently used metrics in machine translation: BLEU (Papineni et al., 2002), ChrF (Popović, 2015), spBLEU (Goyal et al., 2022), BLEURT (Sellam et al., 2020), COMET (Rei et al., 2020).
BLEU and ChrF are n-gram matching heuristics while the rest uses a parametric model to produce a segment-level score of a translation. Comet and CometKiwi are special cases which do not require a reference. We do not include any LLM-based metrics (Fernandes et al., 2023; Kocmi and Federmann, 2023) which are not replicable because of non-publicly available models. Find the specific models, implementation details, and selection rationale in Appendix A.
@@
We focus solely on the pairwise system ranking – deciding which system is better based on a system-level score111Usually the average of all segment level scores. difference between two systems. We refer to this as metric delta ().
Because the order of two systems is arbitrary and do not affect any of our evaluations, only some visualizations, we always specify the first system to be the superior with respect to the human judgement. Therefore human score delta is always positive, while metric delta may be negative, which represents that metric incorrectly ranked the system pair in reverse order to humans.
@@ 3.1 Various Ranges for Metric Deltas
Figure 1 depicts the distribution of system-level score deltas for various metrics. Some metrics have similar ranges, such as ChrF and BLEU, while others use much larger score range (Comet has higher deltas to BLEU) or lower score range (Comet has range of BLEU).
It may be tempting to attempt to bring together these score ranges onto a single scale, e.g. by linear interpolation, perhaps towards BLEU scale. We begin by noting the infeasibility of arriving at a single unified scale. We use the mostly unwritten (and long-debunked (Mathur et al., 2020a)) operating assumption that +1–2 BLEU points denotes a significant finding as an anchor point to illustrate the range of metric deltas on a subset of systems in Figure 2. This figure reports metric deltas for six randomly-selected system pairs from WMT23 data, whose delta was roughly 1 BLEU.
In addition to the wide ranges of scores, we also observe that metrics do not always have the same direction or agreement with human judgment. Reconciling these by projection is not possible, due to an obvious point: metrics differ not just in the range of their scores, but in their accuracies. To better understand the problem, we look next into what are the implications of different thresholds. Specifically, we investigate how different delta correspond to humans being able to differentiate systems.
@@ 3.2 Accuracy of Metric Deltas
Each metric and also humans weights various phenomena differently, especially fluency versus adequacy (Amrhein et al., 2022).
@@
The pairwise accuracy as usually reported (Kocmi et al., 2021; Freitag et al., 2023) represent a value over the full dataset for all system pairs metric deltas. It does not take into consideration the size of the delta between systems, which heavily affects the accuracy; that is, whether the metric gap between two systems was large or small. However, this information is important in establishing equivalency of deltas across metrics.
To investigate this, we use a binning approach on the ToShip23 testset. Pairwise system deltas are sorted, and for each delta level, we group the closest 300 pairs into a same bin. For each bin, we plot the mean delta for that bin against the system-level pairwise accuracy.222Appendix C investigates other sizes of bins than selected 300 system pairs.
@@ 3.3 Aligning Metrics on Accuracy
Practitioners might be interested in getting an intuition behind a particular metric delta, e.g. +0.10 of Comet and how such delta relates to other metrics that they are familiar with. Clearly, the higher the delta, the more likely that human raters would also notice the quality difference between systems. It remains unclear what delta is enough to warrant acceptance. To this end, we use the estimated accuracy results introduced in previous subsection. As the estimated accuracy line is noisy, we fit a curve through the data and use it to derive thresholds for comparing various metric deltas.
@@
The resulting fit is visualized for several metrics in Figure 4. Although not perfect, it provides insight into the metric delta behaviour, specifically comparing different delta in terms of their estimated accuracy. We use the sigmoid functions to calculate estimated accuracy for various levels of delta in Table 2. This is the core result of our work and helps in understanding how different metrics compare to each other.
For example, an improvement of 1.06 BLEU has the same estimated accuracy (65%) as the 0.24 CometKiwi, while 3.35 BLEU has the same estimated accuracy as 0.67 CometKiwi. And +1 improvement on CometKiwi signalize that in >90% scenarios, human annotators would agree with the ranking of CometKiwi, while BLEU never reaches this level of agreement. Note that estimated accuracies are empirical from a given ToShip23 dataset. Therefore, we do not claim that +0.56 Comet yields 80% accuracy for all scenarios but rather that it is as accurate as +2.34 BLEU.
@@ 4.1 Different Domains and Datasets
We derived the thresholds from the ToShip23 dataset. Now, we validate them on WMT data to show how well they transfer. To address the relatively small size of WMT, we first combine the WMT 2022 and 2023 datasets, which yields 1414 system pairs. This dataset contains different set of segment sources and domains, and was evaluated with mix of MQM and DA+SQM human evaluation protocols. In order to test the thresholds, we take scores for all WMT system pairs and convert them into estimated accuracies via devised thresholds. For each estimated accuracy level, we take the closest 300 system pairs and calculate the real accuracy on WMT data. If the mapping would be perfect and we had enough samples, the estimated accuracy would match the real accuracy for each investigated level.
We show the results in Figure 5. In the ideal case, we would expect the real accuracies and estimated accuracies to match, however, the noise from empirical data affects the results. Some metrics are consistently underestimated, such as Comet which has higher real accuracies on WMT dataset that the estimated accuracies. On the other hand, Comet has much lower accuracies on WMT data and our thresholds overestimate it.
Overall, the trend is visible and the thresholds successfully normalize all metrics into a shared space of estimated accuracies. Therefore, we advice to report estimated accuracy when presenting results, together with statistical significance testing and actual metric delta.
@@ 4.2 Language Pair
Notoriously there is a large gap in absolute BLEU scores between languages (Denoual and Lepage, 2005; Post, 2018). This reflects varying properties such as training data size, the attention progress in different language pairs, and target-side morphological complexity.
@@
Figure 6 show the accuracy with only a subset of system pairs depending on a languages. There is some fluctuation between XE and EX, but the behaviour is comparable. This is an interesting finding, since most of the underlying testsets have authentic source (e.g., not using testset in reverse direction, Toral et al., 2018). The CJK group does also perform similarly for CometKiwi, but not for ChrF. This shows that the thresholds are not valid for all metrics in all scenarios and are affected by the point if metrics evaluate all language in the similar fashion or not.
@@ 4.3 Iterated versus Unrelated Systems
Another main difference that affects the evaluation is if the systems are closely related. One point of distinction is between iterated systems (a baseline against specific improvements, produced by the same group) or unrelated system (for example, WMT yearly evaluation which comes from different teams). It has long been known that surface metrics like BLEU work best when evaluating closely related systems (Callison-Burch et al., 2006). It may be easier for both metrics and humans to distinguish improved system over baseline, while comparing unrelated systems adds a difficulty of weighting different styles and errors.
To investigate this, we use the labels of ToShip23 dataset, where some system pairs are improvements over the previous model, while other system pairs are completely unrelated and developed by different teams, similarly to WMT evaluation. Figure 7 confirms the assumption, that unrelated systems are harder to evaluate and that the metric behaves differently. Therefore, automatic metrics are better to rank improved systems which is mostly their use case than unrelated systems.
@@ 4.4 Testset Size
From Figure 8, the metric delta fluctuates but keeps being mostly constant. The variance of the metric delta is higher for small testset sizes (under 500 segments). On the other hand, the p-value associated to the comparison hypothesis goes down simply by having a larger testset. The p-value goes down faster for larger effect sizes.
This is a natural phenomenon of statistical significance testing (Greenland et al., 2016). P-values decrease with an increasing sample size, assuming the null hypothesis does not hold. This is due to the increase in statistical power—the probability that the test correctly rejects the null hypothesis when it is false. Should the null hypothesis hold perfectly, which is rarely the case, increasing the sample size would not systematically affect the p-values. Therefore, it is possible to claim a statistically significant improvement over a baseline model even with a small metric delta, which might not be noticeable by humans, just by using a large-enough testset. This conclusion is not an argument against the use of statistical significance testing, which is important especially when observing smaller deltas.
An interesting research question is if we could fix the testset size to stabilize p-value.
@@ 5.1 Best-performing Metrics
With the ToShip23 dataset, we can also calculate total pairwise accuracy over all system pairs to devise which metric perform the best on up to date the largest dataset of MT human evaluation. We follow the same evaluation as in Table 2 from Kocmi et al. (2021). With twice as large dataset than ToShip21, extended by state-of-the-art systems from 2022 and 2023, we can see how metric perform on system-level ranking. Table 3 shows that the best performing metric over the ToShip23 dataset is CometKiwi by a small margin over xCOMET. CometKiwi is a quality estimation metric, which has an additional bonus of not being affected by reference bias.
@@
Additionally, we can notice the overall accuracy dropped for all metrics in the last two years. This may have several possible explanations and likely does not mean drop in metrics performance:
Different systems: Newer architectures or systems that are closer to each other in performance, thus harder to evaluate by humans
@@
Human bias: The human evaluation protocol changed between years which shifted annotator’s scoring patterns.
However, the absolute pairwise accuracy is not as important as the overall ranking of metrics, as it is heavily affected by the sample of system pairs. We compare to MQM subset of Freitag et al. (2023), which ranks metrics in similar order supporting our findings. There are some notable differences, such as Comet ranking as the worst metric in WMT, while BLEURT is the worst in ToShip23. Since many aspects of the evaluation are different, we do not dive into comparison, but rather highlight the overall picture. ToShip23 corroborates that QE metrics have reached the quality of reference-based metrics, as well as the (already well-established) fact that lexical-based metrics are not as useful for evaluating high-resource MT models these days.

### v2

@@ Abstract
Ten years ago, a single metric, BLEU, governed progress in machine translation research. For better or worse, there is no such consensus today, and consequently it is difficult for researchers to develop and retain intuitions about metric deltas that drove earlier research and deployment decisions. This paper investigates the “dynamic range” of a number of modern metrics in an effort to provide a collective understanding of the meaning of differences in scores both within and among metrics; in other words, we ask what point difference in metric is required between two systems for humans to notice? We conduct our evaluation on a new large dataset, ToShip23, using it to discover deltas at which metrics achieve system-level differences that are meaningful to humans, which we measure by pairwise system accuracy. We additionally show that this method of establishing delta-accuracy is more stable than the standard use of statistical p-values in regards to testset size. Where data size permits, we also explore the effect of metric deltas and accuracy across finer-grained features such as translation direction, domain, and system closeness.
Navigating the Metrics Maze:
Reconciling Score Magnitudes and Accuracies
Tom Kocmi1 Vilém Zouhar2 Christian Federmann1 Matt Post1 1Microsoft 2ETH Zürich {tomkocmi,chrife,mattpost}@microsoft.com vzouhar@ethz.ch
Online tool: kocmitom.github.io/MT-Thresholds
@@ 1 Introduction
A decade ago, the BLEU metric served as the default metric for machine translation evaluation. It was not without its criticisms (Hovy and Ravichandran, 2003; Callison-Burch et al., 2006; Belz and Reiter, 2006) or compelling alternatives (Banerjee and Lavie, 2005; Popović, 2015), but a combination of adequate performance, robustness to new languages, simplicity, understandability—and also inertia—helped it retain this position. This is no longer the case. BLEU’s deficiencies quickly became apparent as deep learning approaches to machine translation replaced the earlier symbolic paradigms (Mathur et al., 2020a). Today, a number of metrics—themselves deep-learning based—compete in an ecosystem where there is no longer any dominant, default metric.
This situation creates a problem for researchers working to keep abreast of developments in the field. Different metrics, including different models within the same metric family, have different dynamic ranges, i.e., the range of scores one can expect to see. Furthermore, the metric delta, i.e., the score difference signifying a meaningful change in performance between two systems, also varies across metrics. It is perhaps understandable that some practitioners therefore continue to use BLEU, as well, if only to ground their understanding.
@@
In this work we:
§3.2
Empirically investigate the estimated accuracy for multiple metrics, human ability to perceive quality difference;
§3.3
Provide thresholds for popular metrics to help reviewers and practitioners interpret results;
§4.1
Validate our estimated accuracies on WMT testsets and §4.2 investigate the effect of different language groups;
§4.3
Show that string-based metrics, such as BLEU, should never be used to evaluate unrelated systems;
§4.4
Show that statistical significance testing is insufficient to determine model improvement especially as it is affected by the testset size, but is important for small deltas;
§5.1
Assess quality of automatic metrics over 6530 system pairs;
§5.2
Summarize recommendations for machine translation evaluation.
@@ 2 Experimental Setup
We perform experiments related to evaluation of MT outputs based on a proprietary dataset ToShip23 which is of a magnitude larger than any publicly available data and enables more fine grained glimpse into the metrics behaviour. The dataset is an extended version of ToShip21 dataset (Kocmi et al., 2021) with details described in Appendix B. We also use data from the annual WMT evaluation campaigns to validate our results, specifically the metrics shared task (Freitag et al., 2022b, 2023), to make results replicable. We only use MQM (Freitag et al., 2021a) and DA+SQM (Kocmi et al., 2022) subset of human evaluated systems because reference-based DA (Bojar et al., 2016) is suboptimal for the evaluation of modern MT systems (Freitag et al., 2022b). See Table 1 for an overview of dataset sizes.
@@
We evaluate the most frequently used metrics in machine translation: BLEU (Papineni et al., 2002), ChrF (Popović, 2015), spBLEU (Goyal et al., 2022), BLEURT (Sellam et al., 2020), COMET (Rei et al., 2020). BLEU and ChrF are n-gram matching heuristics while the rest uses a parametric model to produce a segment-level score of a translation. Comet and CometKiwi are special cases which do not require a reference. We do not include any LLM-based metrics (Fernandes et al., 2023; Kocmi and Federmann, 2023) which are not replicable because of non-publicly available models. Find the specific models, implementation details, and our selection rationale in Appendix A.
@@
We focus solely on the pairwise system ranking: deciding which system is better based on a system-level score (usually average of all segment-level scores) difference between two systems. We refer to this as metric delta ().
@@ 3.1 Various Ranges for Metric Deltas
Figure 1 depicts the distribution of system-level score deltas for various metrics. Some metrics have similar ranges, such as ChrF and BLEU, while others use a much larger score range (Comet has higher deltas to BLEU) or lower score range (Comet has range of BLEU).
In addition to the wide ranges of scores, we also observe that metrics do not always have the same direction or agreement with human judgment, which results in their different performance as measured via accuracy (see Appendix C for more details).
It may be tempting to attempt to bring together these score ranges onto a single scale, say by linear interpolation, perhaps towards BLEU scale. But reconciling metrics by projection is not possible, due to an obvious point: metrics differ not just in the range of their scores, but in their accuracies. To better understand the problem, we look next into what are the implications of different levels of metric deltas. Specifically, we investigate how different delta correspond to humans being able to differentiate systems.
@@ 3.2 Accuracy of Metric Deltas
The pairwise accuracy as usually reported (Kocmi et al., 2021; Freitag et al., 2023) represents a value over the full dataset for all system-pair metric deltas. It does not take into consideration the size of the delta between systems, which heavily affects the accuracy; that is, whether the metric gap between two systems was large or small. However, this information is important in establishing equivalency of deltas across metrics.
To investigate this, we use a binning approach on the ToShip23 testset. Pairwise system deltas are sorted, and for each delta level, we group the closest 300 pairs into a same bin. For each bin, we plot the mean delta for that bin against the system-level pairwise accuracy.111Appendix D investigates other sizes of bins than selected 300 system pairs.
@@ 3.3 Aligning Metrics on Accuracy
Practitioners might be interested in getting an intuition behind a particular metric delta, e.g., +0.10 of Comet and how such delta relates to other metrics that they are familiar with. Clearly, the higher the delta, the more likely that human raters would also notice the quality difference between systems. It remains unclear what delta is enough to warrant acceptance. To this end, we use the estimated accuracy results introduced in previous subsection. As the estimated accuracy line is noisy, we fit a curve through the data and use it to derive thresholds for comparing various metric deltas.
@@
The resulting fit is visualized in Figure 3. Although not perfect, it offers insight into the metric delta behaviour, specifically comparing different different deltas’ estimated accuracy. We use the sigmoid functions to calculate estimated accuracy for various levels of delta in Table 2. This is the core result of our work and helps in understanding how different metrics compare to each other.
For example, an improvement of 1.06 BLEU has the same estimated accuracy (65%) as the 0.24 CometKiwi, while 3.35 BLEU has the same estimated accuracy as 0.67 CometKiwi. And +1 improvement on CometKiwi signals that in >90% scenarios, human annotators would agree with the ranking of CometKiwi, while BLEU never reaches this level of agreement. Note that estimated accuracies are empirical from a given ToShip23 dataset. Therefore, we do not claim that +0.56 Comet yields 80% accuracy for all scenarios but rather that it is as accurate as +2.34 BLEU.
@@ 4.1 Different Domains and Datasets
We derived the thresholds from ToShip23. Now, we validate them on WMT data to show how well they transfer. To address the relatively small size of WMT, we first combine the WMT 2022 and 2023 datasets, yielding 1414 system pairs. This dataset contains different set of segment sources and domains, and was evaluated with mix of MQM and DA+SQM human evaluation protocols. In order to test the thresholds, we take scores for all WMT system pairs and convert them into estimated accuracies via devised thresholds. For each estimated accuracy level, we take the closest 300 system pairs and calculate the real accuracy on WMT data. If the mapping would be perfect and we had enough samples, the estimated accuracy would match the real accuracy for each investigated level.
We show the results in Figure 4. In the ideal case, we would expect the real accuracies and estimated accuracies to match; however, the noise from empirical data affects the results. Some metrics are consistently underestimated, such as Comet, which has higher real accuracies on WMT dataset that the estimated accuracies. On the other hand, Comet has much lower accuracies on WMT data and our thresholds overestimate it.
Overall, the trend is clear and the thresholds normalize all metrics into a shared space of estimated accuracies. Therefore, we advise reporting accuracy when presenting results, together with significance testing and metric delta.
@@ 4.2 Language Pair
Notoriously, a large gap in absolute BLEU scores exists between languages (Denoual and Lepage, 2005; Post, 2018). This reflects properties like data sizes, attention progress in different languages, and target-side morphological complexity.
@@
Figure 5 show the accuracy with a subset of system pairs depending on a languages. There is some fluctuation between XE and EX, but the behaviour is comparable. This is interesting, since most of the underlying testsets have authentic source (e.g., not using testset in reverse direction, Toral et al., 2018). The CJK group does also perform similarly for CometKiwi, but not for ChrF. This shows the thresholds are invalid for all metrics and scenarios and are affected by whether metrics evaluate all language similarly or not.
@@ 4.3 Iterated versus Unrelated Systems
Another main difference that affects the evaluation is if the systems are closely related. Key point of distinction is between iterated systems (a baseline system against specific improvements, produced by the same research group) or unrelated system (for example, WMT yearly evaluation which comes from different teams and systems produce vastly different translations). It has long been known that surface metrics like BLEU work best when evaluating closely-related iterated systems (Callison-Burch et al., 2006). It may be easier for both metrics and humans to distinguish an iterated system over its baseline, because comparing unrelated systems adds a difficulty of weighting different styles and errors.
To investigate this, we use the system labels of ToShip23 dataset, where some system pairs are baseline model and it’s improved iterated model, while other system pairs are completely unrelated and developed by different teams, similarly to WMT evaluation. Figure 6 confirms the assumption that unrelated systems are much harder to evaluate and that the metric behaves differently. Therefore, automatic metrics are better to rank iterated systems than unrelated systems. While pretrained metrics, such as CometKiwi, seems to be robust enough for comparing both types of system pairs, other metrics such as BLEU have much harder time to distinguish unrelated systems. This effect should be investigated to larger detail in future work.
For example, +2 BLEU on iterated models has an accuracy with humans of about 90%, the same +2 BLEU on unrelated systems are barely better than toss of a coin (55%). This shows, that some metrics (specifically BLEU, ChrF, spBLEU) should not be used to evaluate unrelated systems. This findings was also suggested by Berg-Kirkpatrick et al. (2012), who showed that you need to get about one third larger BLEU improvements for unrelated systems to reach the same p-value.
Therefore, string-based metrics, such as BLEU, ChrF, or spBLEU, should never be used to compare unrelated systems.
@@ 4.4 Testset Size
From Figure 7, the metric delta fluctuates but keeps being mostly constant. The variance of the metric delta is higher for small testset sizes (under 500 segments). On the other hand, the p-value associated to the comparison hypothesis goes down simply by having a larger testset, phenomena shown for MT by Berg-Kirkpatrick et al. (2012).
This is a natural phenomenon of statistical significance testing (Greenland et al., 2016). P-values decrease with an increasing sample size, assuming the null hypothesis does not hold. This is due to the increase in statistical power—the probability that the test correctly rejects the null hypothesis when it is false. Should the null hypothesis hold perfectly, which is rarely the case, increasing the sample size would not systematically affect the p-values. Therefore, it is possible to claim a statistically significant improvement over a baseline model even with a small metric delta, which might not be noticeable by humans, just by using a large-enough testset. This conclusion is not an argument against the use of statistical significance testing, which remains important, especially when observing smaller deltas.
@@ 5.1 Best-performing Metrics
With the ToShip23 dataset, we can also calculate total pairwise accuracy over all system pairs to devise which metrics perform the best on the (to date) largest dataset of MT human evaluation. We follow the same evaluation as in Table 2 from Kocmi et al. (2021). Twice as large dataset than ToShip21, extended by state-of-the-art systems from 2022 and 2023, we can see how metrics perform on system-level rankings. Table 3 shows that the best performing metric over the ToShip23 dataset is CometKiwi by a small margin over xCOMET. CometKiwi is a quality estimation metric, which has an additional bonus of not being affected by reference bias.
@@
Additionally, we notice the overall accuracy dropped for all metrics in the last two years. This does not necessarily signify a drop in metric performance, but may have several other explanations:
Different systems: Newer architectures or systems are closer to each other in performance, thus harder to evaluate by humans
@@
Human bias: The evaluation protocol changed, which may have shifted annotator’s scoring patterns.
However, the absolute pairwise accuracy is less important than the ranking of metrics, as it is heavily affected by the system pairs. We compare to MQM subset of Freitag et al. (2023), which ranks metrics in similar order supporting our findings. There are some notable differences, such as Comet ranking as the worst metric in WMT, while BLEURT is the worst in ToShip23. Since many aspects of the evaluation are different, we do not dive into a comparison, but rather highlight the overall picture. ToShip23 corroborates that QE metrics have reached the quality of reference-based metrics, as well as the (already well-established) fact that lexical-based metrics are not useful for evaluating high-resource MT models these days.